Calculate the mass of iron(|||) oxide (rust) produced by the reaction of 500 g of iron with oxygen from the air
1
Expert's answer
2020-05-13T14:20:12-0400

4Fe + 3O2 = 2Fe2O3

"Moles =mass\/rmm"

If 500g of Fe react then, 500/55.8=9moles reacted.

The mole ratio of Fe:Fe2O3=4:2= 2:1

Therefore if 9 moles of Fe reacted 4.5 moles of Fe2O3 are produced

"Mass=moles \times rmm"

Mass of Fe2O3= 4.5* 159.7= 718.7g